OpenStudy (anonymous):

perimeter of rectangle is 40ft. length is 8ft longer then width. Find dimensions let x be length and y width 2x+2y=

OpenStudy (anonymous):

2x+2y=40 l=w+8 2(w+8)+2w=40 2w+16+2w=40 4w=24 w=6

OpenStudy (anonymous):

switch all the w's to y

OpenStudy (anonymous):

Here: \[ 2l+2w=40\\ l=8+w \]Substituting: \[ 2(8+w)+2w=40\\ 16+4w=40\\ 4w=24\\ w=6\\ l=8+6=14 \]

OpenStudy (anonymous):

thanks